# Statistical inference on partially shape-constrained function-on-scalar linear regression models

### v1

@@ 2.2 Estimation: Shape-constrained kernel-weighted least squares
For each $t\in[0,1]$, let $\mathbb{B}_{0}(t)=\big{[}\boldsymbol{\beta}_{0}(t),\dot{\boldsymbol{\beta}}_{0}(t)\big{]}$ be the $p\times 2$ matrix, where $\boldsymbol{\beta}_{0}(t)$ is the true vector coefficient function such that $\ddot{\boldsymbol{\beta}}_{0}(t)=\big{(}\beta_{1}^{\prime\prime}(t),\ldots,\beta_{p}^{\prime\prime}(t)\big{)}^{\top}$ exists and is continuous. Suppose conditions $(K0)$$(K4)$ in Supplementray Material S.1 hold. Then, we have
$\sup_{t\in[0,1]}\big{\|}\widetilde{\boldsymbol{\beta}}(t)-\boldsymbol{\beta}_{0}(t)\big{\|}=O_{P}\bigg{(}h^{2}+\sqrt{\frac{\log n}{nh}}\bigg{)}.$ (2.5)
The implication of conditions $(K0)$$(K4)$ are provided in Supplementary Material S.1. It is worth mentioning that Proposition 2.2.1 is valid even if only a few longitudinal observations are available for some subjects. For example, it is common in observational studies that dense observations of functional responses for some subjects may not be available regardless of the sample size $n$, i.e., the minimum number of longitudinal observations $\lambda_{n}=\min_{1\leq i\leq n}L_{i}$ is bounded. Therefore, the technical condition on $\lambda_{n}$ in $(K4)$ is not a restriction because we do not require $\lambda_{n}\to\infty$ as $n\to\infty$.
@@
For technical conditions $(K0)$$(K4)$, $k>2.5$ and $a=1/5$ are commonly adopted in the standard theory of kernel smoothing [10]. Then, Theorem 2.2.1 gives $\sup_{t\in[0,1]}\|\tilde{\boldsymbol{\beta}}(t)-\boldsymbol{\beta}_{0}(t)\|=O_{P}(n^{-2/5}\sqrt{\log n})$. It follows that $\int_{0}^{1}\big{\|}\widetilde{\boldsymbol{\beta}}(t)-\boldsymbol{\beta}_{0}(t)\big{\|}^{2}\,\mathrm{d}t=O_{P}(n^{-2/5}\sqrt{\log n})$.
For the shape-constrained estimation of $\widehat{\boldsymbol{\beta}}$ under $H_{0}$, we extend the kernel-weighted least squares for nonparametric regression with scalar responses [34], where side conditions are empirically examined over a fine grid $\mathcal{G}_{M}=\{t_{m}\in[0,1]:m=0,1,\ldots,M\}$ with $10=t_{0}<t_{1}<\cdots<t_{M-1}<t_{M}=$ 1 so that $\hat{\beta}_{j}$ satisfies the null hypothesis (2.2) over $\mathcal{G}_{j}=\mathcal{G}|_{\mathcal{I}_{j}}=\{t_{j_{m}}\in\mathcal{I}_{j}:m=0,1,\ldots,M_{j}\}$ for each $j\in J$. Specifically, if we consider “$H_{0,j}:\textrm{$\beta_{j}$ is monotone increasing on $\mathcal{I}_{j}$}$,” we propose to substitute
@@
Similarly, we empirically validate the partial hypothesis “$H_{0,j}:\textrm{$\beta_{k}$ is convex on $\mathcal{I}_{j}$}$” as
$H_{0,j}^{\prime}:\frac{\beta_{j}(t_{j_{m}})-\beta_{j}(t_{j_{m-1}})}{t_{j_{m}}-t_{j_{m-1}}}\geq\beta_{j}^{\prime}(t_{j_{m-1}})\,\,\,\textrm{forall}\,\,\,m=1,\ldots,M_{j}.$ (2.7)
To define the shape-constrained kernel-weighted least squares for the function-on-scalar regression model (2.1) under the empirical null hypothesis on the grid $\mathcal{G}_{M}$, we note that
$\begin{split}\mathcal{L}_{n}\big{(}\mathbb{B}(t)\big{)}&=\sum_{i=1}^{n}\frac{1}{L_{i}}\sum_{\ell=1}^{L_{i}}\Big{(}\mathbb{W}_{i,\ell}(t)^{\top}\big{[}\mathrm{vec}\big{(}\mathbb{B}(t)\big{)}-\mathrm{vec}\big{(}\widetilde{\mathbb{B}}(t)\big{)}\big{]}-\tilde{e}_{i,\ell}(t)\Big{)}^{2}K_{h}(T_{i,\ell}-t)\\ &=\mathcal{Q}_{n}\big{(}\mathbb{B}(t),\widetilde{\mathbb{B}}(t)\big{)}+\frac{1}{n}\sum_{i=1}^{n}\frac{1}{L_{i}}\sum_{\ell=1}^{L_{i}}K_{h}(T_{i,\ell}-t)\tilde{e}_{i,\ell}(t)^{2},\end{split}$ (2.8)
where $\tilde{e}_{i,\ell}=Y_{i,\ell}-\mathbb{W}_{i,\ell}^{\top}\mathrm{vec}\big{(}\widetilde{\mathbb{B}}(t)\big{)}$ is the unconstrained residual of fitting $Y_{i,\ell}$ and $\mathcal{Q}_{n}\big{(}\mathbb{B}(t),\widetilde{\mathbb{B}}(t)\big{)}=\mathrm{vec}\big{(}\mathbb{B}(t)-\widetilde{\mathbb{B}}(t)\big{)}^{\top}\widehat{\Psi}(t)\mathrm{vec}\big{(}\mathbb{B}(t)-\widetilde{\mathbb{B}}(t)\big{)}$ with $\widehat{\Psi}(t)=n^{-1}\sum_{i=1}^{n}L_{i}^{-1}\sum_{\ell=1}^{L_{i}}\mathbb{W}_{i,\ell}(t)\mathbb{W}_{i,\ell}(t)^{\top}K_{h}(T_{i,\ell}-t)$. To get (2.8), we used the fact that the unconstrained estimator $\widetilde{\mathbb{B}}(t)$ solves the estimating equation $\partial\mathcal{L}_{n}\big{(}\mathbb{B}(t)\big{)}/\partial\mathrm{vec}\big{(}\mathbb{B}(t)\big{)}=0_{2p}$, or equivalently $n^{-1}\sum_{i=1}^{n}L_{i}^{-1}\sum_{\ell=1}^{L_{i}}\mathbb{W}_{i,\ell}(t)\tilde{e}_{i,\ell}K_{h}(T_{i,\ell}-t)=0_{2p}$, where $0_{2p}=(0,\ldots,0)^{\top}$ denotes the zero vector in $\mathbb{R}^{2p}$. Noting that the minimization of $\mathcal{L}_{n}\big{(}\mathbb{B}(t)\big{)}$ only depends on the first term of (2.8) given the unconstrained estimator $\widetilde{B}(t)$, we propose the constrained optimization problem as follows.
where $0_{k}\in\mathbb{R}^{k}$ and $O_{k\times\ell}\in\mathbb{R}^{k\times\ell}$ are the $k$-vector and $(k,\ell)$-matrix of zeros, respectively, and $A_{j}\in\mathbb{R}^{M_{j}\times 2p(M+1)}$ is the linear constraint matrix associated with the individual hypothesis $H_{0,j}$. For the monotone increasing hypothesis (2.6), we set $A_{j}=A_{\mathrm{mono}}(\mathcal{I}_{j})\otimes\big{[}\mathrm{e}_{j}^{\top},0_{p}^{\top}\big{]}$, where $\textrm{e}_{j}\in\mathbb{R}^{p}$ is the unit vector with $1$ only at its $j$-th coordinate and $A_{\mathrm{mono}}(\mathcal{I}_{j})\in\mathbb{R}^{M_{j}\times(M_{j}+1)}$ is defined as
@@
Similarly, for the convexity hypothesis (2.7), we set we set $A_{j}=A_{\mathrm{conv},0}(\mathcal{I}_{j})\otimes\big{[}\mathrm{e}_{j}^{\top},0_{p}^{\top}\big{]}-A_{\mathrm{conv},1}(\mathcal{I}_{j})\otimes\big{[}0_{p}^{\top},\mathrm{e}_{j}^{\top}\big{]}$, where $A_{\mathrm{conv},0}(\mathcal{I}_{k}),A_{\mathrm{conv},1}(\mathcal{I}_{j})\in\mathbb{R}^{M_{j}\times(M_{j}+1)}$ are defined as
$\begin{split}A_{\mathrm{conv},0}(\mathcal{I}_{j})&=\left[\begin{array}[]{ccc c cc}\frac{-1}{t_{j_{1}}-t_{j_{0}}}&\frac{1}{t_{j_{1}}-t_{j_{0}}}&0&\cdots&0&0\\ 0&\frac{-1}{t_{j_{2}}-t_{j_{1}}}&\frac{1}{t_{j_{2}}-t_{j_{1}}}&\cdots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&0&\cdots&\frac{-1}{t_{j_{M_{j}}}-t_{j_{M_{j}-1}}}&\frac{1}{t_{j_{M_{j}}}-t_{j_{M_{j}-1}}}\end{array}\right]\end{split}$ (2.15)
and $A_{\mathrm{conv},1}(\mathcal{I}_{j})=\big{[}I_{M_{j}},0_{M_{j}}\big{]}$. This illustrates that the constrained estimates $\widehat{\mathbb{B}}(t_{0}),\ldots,\widehat{\mathbb{B}}(t_{M})$ can be obtained by the standard quadratic programming with $\sum_{j\in J}M_{j}$ linear constraints.
@@
Suppose we hypothesize “$H_{0,j}:\textrm{$\beta_{j}$ is monotone increasing on $\mathcal{I}_{j,1}$ and convex on $\mathcal{I}_{j,2}$}$,” where $\mathcal{I}_{j,1}$ and $\mathcal{I}_{j,2}$ may or may not overlap. Then, the side conditions corresponding to the null hypothesis can also be examined by imposing both linear constraint matrices $A_{\mathrm{mono}}(\mathcal{I}_{j,1})\otimes\big{[}\mathrm{e}_{j}^{\top},0_{p}^{\top}\big{]}$ and $A_{\mathrm{conv},0}(\mathcal{I}_{j,2})\otimes\big{[}\mathrm{e}_{j}^{\top},0_{p}^{\top}\big{]}-A_{\mathrm{conv},1}(\mathcal{I}_{j,2})\otimes\big{[}0_{p}^{\top},\mathrm{e}_{j}^{\top}\big{]}$.
@@
Suppose the conditions in Proposition 2.2.1 hold. For any $M\geq 1$, let $\mathcal{G}_{M}=\{t_{m}\in[0,1]:m=0,1,\ldots,M\}$ be the finite grid used in (2.9) with $10=t_{0}<t_{1}<\cdots<t_{M}=$ 1. Under the null hypothesis $H_{0}$ in (2.2), we have
$\max_{t\in\mathcal{G}_{M}}\big{\|}\widehat{\boldsymbol{\beta}}(t)-\boldsymbol{\beta}_{0}(t)\big{\|}=O_{P}\bigg{(}h^{2}+\sqrt{\frac{\log n}{nh}}\bigg{)}.$ (2.16)
Theorem 2.2.2 indicates that the shape-constrained kernel least square estimator achieves the same rate of convergence as the unconstrained estimator under the null hypothesis. Moreover, suppose the empirical distribution of $\mathcal{G}_{M}$ converges to the uniform distribution on $[0,1]$. Then, it can be verified that $\int_{0}^{1}\big{\|}\widehat{\boldsymbol{\beta}}(t)-\boldsymbol{\beta}_{0}(t)\big{\|}\,\mathrm{d}t=O_{P}(n^{-2/5}\sqrt{\log n})$ similarly to Remark 1.

### v2

@@ 5. $(K4)$
For each $t\in[0,1]$, let $\mathbb{B}(t)=\big{[}\boldsymbol{\beta}(t),\dot{\boldsymbol{\beta}}(t)\big{]}$ be the $p\times 2$ matrix, where $\ddot{\boldsymbol{\beta}}(t)=\big{(}\beta_{1}^{\prime\prime}(t),\ldots,\beta_{p}^{\prime\prime}(t)\big{)}^{\top}$ exists and is continuous. Suppose conditions $(K0)$$(K4)$ hold. Then, we have
$\sup_{t\in[0,1]}\|\widetilde{\boldsymbol{\beta}}(t)-\boldsymbol{\beta}(t)\|=O_{P}\bigg{(}h^{2}+\sqrt{\frac{\log n}{nh}}\bigg{)}.$ (2.2)
Proposition 2.2.1 establishes the large-sample properties of the unconstrained estimator. While the presented rate of convergence aligns with standard results for kernel smoothing in the functional data analysis literature, we have relaxed certain conditions to enhance flexibility for applications to longitudinal data, which are more common in practice, as illustrated in the Real Data Applications section. Specifically, Proposition 2.2.1 is valid even if only a few longitudinal observations are available for some subjects. For example, it is common in observational studies that dense observations of functional responses for some subjects may not be available regardless of the sample size $n$, i.e., the minimum number of longitudinal observations $\lambda_{n}=\min_{1\leq i\leq n}L_{i}$ is bounded. Therefore, the technical condition on $\lambda_{n}$ in $(K4)$ is not a restriction because we do not require $\lambda_{n}\to\infty$ as $n\to\infty$. Proposition 2.2.1 also serves as a fundamental component in establishing one of our main results, which is presented in Theorem 2.2.2.
@@
For technical conditions $(K0)$$(K4)$, $k>2.5$ and $a=1/5$ are commonly adopted in the standard theory of kernel smoothing [10]. Then, Proposition 2.2.1 gives $\sup_{t\in[0,1]}\|\tilde{\boldsymbol{\beta}}(t)-\boldsymbol{\beta}(t)\|=O_{P}(n^{-2/5}\sqrt{\log n})$. It follows that $\int_{0}^{1}\|\widetilde{\boldsymbol{\beta}}(t)-\boldsymbol{\beta}(t)\|\,\mathrm{d}t=O_{P}(n^{-2/5}\sqrt{\log n})$.
For the shape-constrained estimation of $\boldsymbol{\beta}$, we propose extending the shape-constrained kernel-weighted least squares (SCKLS) approach to function-on-scalar linear regression (FoSR). [42] introduced the shape-constrained kernel-weighted least squares (SCKLS) method for nonparametric regression, where the conditional expectation of a scalar response given multiple covariates is subject to global shape constraints. Our proposed method preserves the nonparametric flexibility of the functional regression model by leveraging kernel estimation while focusing on partial shape constraints applied to individual regression coefficient functions, in contrast to the global shape constraints imposed on the multivariate regression function in [42]. To this end, we note that
$\begin{split}\mathcal{L}_{n}\big{(}\mathbb{B}(t)\big{)}&=\frac{1}{n}\sum_{i=1}^{n}\frac{1}{L_{i}}\sum_{\ell=1}^{L_{i}}\Big{(}\mathbb{W}_{i,\ell}(t)^{\top}\big{[}\mathrm{vec}\big{(}\mathbb{B}(t)\big{)}-\mathrm{vec}\big{(}\widetilde{\mathbb{B}}(t)\big{)}\big{]}-\tilde{e}_{i,\ell}(t)\Big{)}^{2}K_{h}(T_{i,\ell}-t)\\ &=\mathcal{Q}_{n}\big{(}\mathbb{B}(t),\widetilde{\mathbb{B}}(t)\big{)}+\frac{1}{n}\sum_{i=1}^{n}\frac{1}{L_{i}}\sum_{\ell=1}^{L_{i}}K_{h}(T_{i,\ell}-t)\tilde{e}_{i,\ell}(t)^{2},\end{split}$ (2.3)
where $\tilde{e}_{i,\ell}=Y_{i,\ell}-\mathbb{W}_{i,\ell}^{\top}\mathrm{vec}\big{(}\widetilde{\mathbb{B}}(t)\big{)}$ is the unconstrained residual of fitting $Y_{i,\ell}$ and $\mathcal{Q}_{n}\big{(}\mathbb{B}(t),\widetilde{\mathbb{B}}(t)\big{)}=\mathrm{vec}\big{(}\mathbb{B}(t)-\widetilde{\mathbb{B}}(t)\big{)}^{\top}\widehat{\Psi}(t)\mathrm{vec}\big{(}\mathbb{B}(t)-\widetilde{\mathbb{B}}(t)\big{)}$ with $\widehat{\Psi}(t)=n^{-1}\sum_{i=1}^{n}L_{i}^{-1}\sum_{\ell=1}^{L_{i}}\mathbb{W}_{i,\ell}(t)\mathbb{W}_{i,\ell}(t)^{\top}K_{h}(T_{i,\ell}-t)$. To get (2.3), we used the fact that the unconstrained estimator $\widetilde{\mathbb{B}}(t)$ solves the estimating equation $\partial\mathcal{L}_{n}\big{(}\mathbb{B}(t)\big{)}/\partial\mathrm{vec}\big{(}\mathbb{B}(t)\big{)}=0_{2p}$, or equivalently $n^{-1}\sum_{i=1}^{n}L_{i}^{-1}\sum_{\ell=1}^{L_{i}}\mathbb{W}_{i,\ell}(t)\tilde{e}_{i,\ell}K_{h}(T_{i,\ell}-t)=0_{2p}$, where $0_{2p}=(0,\ldots,0)^{\top}$ denotes the zero vector in $\mathbb{R}^{2p}$.
We note that the minimization of $\mathcal{L}_{n}\big{(}\mathbb{B}(t)\big{)}$ only depends on the first term of (2.3) given the unconstrained estimator $\widetilde{B}(t)$. Based on this observation, we empirically examine shape constraints as side conditions over a fine grid $\mathcal{G}_{M}=\{t_{m}\in[0,1]:m=0,1,\ldots,M\}$ with $10=t_{0}<t_{1}<\cdots<t_{M-1}<t_{M}=$ 1 so that the resulting estimator of each individual regression coefficient fujction estimator $\hat{\beta}_{j}$ satisfies the null hypothesis (2.20) over $\mathcal{G}_{j}=\mathcal{G}|_{\mathcal{I}_{j}}=\{t_{j_{m}}\in\mathcal{I}_{j}:m=0,1,\ldots,M_{j}\}$ for each $j\in J$. Specifically, we consider the following constrained minimization problem:
where $0_{k}\in\mathbb{R}^{k}$ and $O_{k\times\ell}\in\mathbb{R}^{k\times\ell}$ are the $k$-vector and $(k,\ell)$-matrix of zeros, respectively, and $A_{j}\in\mathbb{R}^{M_{j}\times 2p(M+1)}$ is the linear constraint matrix associated with the individual hypothesis $H_{0,j}$. Below, we have collected representative examples illustrating how the partial shape constraints can be translated into linear inequality constraints.
Partial monotonicity: If we consider $H_{0,j}$: $\beta_{j}$ is partially monotone increasing on $\mathcal{I}_{j}$, we propose to substitute
@@
Then, we set $A_{j}=A_{\mathrm{mono}}(\mathcal{I}_{j})\otimes\big{[}\mathrm{e}_{j}^{\top},0_{p}^{\top}\big{]}$, where $\textrm{e}_{j}\in\mathbb{R}^{p}$ is the unit vector with $1$ only at its $j$-th coordinate and $A_{\mathrm{mono}}(\mathcal{I}_{j})\in\mathbb{R}^{M_{j}\times(M_{j}+1)}$ is defined as
@@
Partial convexity: If we want to validate the partial hypothesis $H_{0,j}$: $\beta_{j}$ is partially convex on $\mathcal{I}_{j}$ as
$H_{0,j}^{\prime}:\frac{\beta_{j}(t_{j_{m}})-\beta_{j}(t_{j_{m-1}})}{t_{j_{m}}-t_{j_{m-1}}}\geq\beta_{j}^{\prime}(t_{j_{m-1}})\,\,\,\textrm{forall}\,\,\,m=1,\ldots,M_{j},$ (2.11)
we set $A_{j}=A_{\mathrm{conv},0}(\mathcal{I}_{j})\otimes\big{[}\mathrm{e}_{j}^{\top},0_{p}^{\top}\big{]}-A_{\mathrm{conv},1}(\mathcal{I}_{j})\otimes\big{[}0_{p}^{\top},\mathrm{e}_{j}^{\top}\big{]}$, where $M_{j}\times(M_{j}+1)$ matrices $A_{\mathrm{conv},0}(\mathcal{I}_{k})$ and $A_{\mathrm{conv},1}(\mathcal{I}_{j})$ are defined as
$\begin{split}A_{\mathrm{conv},0}(\mathcal{I}_{j})&=\left[\begin{array}[]{ccc c cc}\frac{-1}{t_{j_{1}}-t_{j_{0}}}&\frac{1}{t_{j_{1}}-t_{j_{0}}}&0&\cdots&0&0\\ 0&\frac{-1}{t_{j_{2}}-t_{j_{1}}}&\frac{1}{t_{j_{2}}-t_{j_{1}}}&\cdots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&0&\cdots&\frac{-1}{t_{j_{M_{j}}}-t_{j_{M_{j}-1}}}&\frac{1}{t_{j_{M_{j}}}-t_{j_{M_{j}-1}}}\end{array}\right],\end{split}$ (2.12)
and $A_{\mathrm{conv},1}(\mathcal{I}_{j})=\big{[}I_{M_{j}},0_{M_{j}}\big{]}$.
Partial positivity: If we empirically test the partial hypothesis $H_{0,j}$: $\beta_{j}$ is partially positive on $\mathcal{I}_{j}$ as
$H_{0,j}^{\prime}:\beta_{j}(t_{j_{m}})\geq 0\,\,\,\textrm{for all}\,\,\,m=0,1,\ldots,M_{j},$ (2.13)
and we set $A_{j}=A_{\mathrm{postive}}(\mathcal{I}_{j})\otimes\big{[}\mathrm{e}_{j}^{\top},0_{p}^{\top}\big{]}$, where $A_{\mathrm{positive}}(\mathcal{I}_{j})$ is the identity matrix in $\mathbb{R}^{(M_{j}+1)\times(M_{j}+1)}$.
@@
The above examples illustrate that the constrained estimates $\widehat{\mathbb{B}}(t)$ can be obtained by the standard quadratic programming with $\sum_{j\in J}M_{j}$ linear constraints, and one can extend our framework to general shape constraints, including partial boundedness and combinations of the above options. For example, suppose we hypothesize $H_{0,j}$: $\beta_{j}$ is monotone increasing on $\mathcal{I}_{j,1}$ and convex on $\mathcal{I}_{j,2}$, where $\mathcal{I}_{j,1}$ and $\mathcal{I}_{j,2}$ may or may not overlap. Then, the side conditions corresponding to the null hypothesis can also be examined by imposing both linear constraint matrices $A_{\mathrm{mono}}(\mathcal{I}_{j,1})\otimes\big{[}\mathrm{e}_{j}^{\top},0_{p}^{\top}\big{]}$ and $A_{\mathrm{conv},0}(\mathcal{I}_{j,2})\otimes\big{[}\mathrm{e}_{j}^{\top},0_{p}^{\top}\big{]}-A_{\mathrm{conv},1}(\mathcal{I}_{j,2})\otimes\big{[}0_{p}^{\top},\mathrm{e}_{j}^{\top}\big{]}$.
We note that the objective function in (2.4) can be expressed as a quadratic form in terms of the $2p(M+1)$-dimensional vector $\big{(}\mathrm{vec}(\mathbb{B}(t_{m})-\widetilde{\mathbb{B}}(t_{m})):m=0,1,\ldots,M\big{)}$, where the associated Gram matrix is a block diagonal matrix composed of $\widehat{\Psi}(t_{0}),\widehat{\Psi}(t_{1}),\ldots,\widehat{\Psi}(t_{M})$. To ensure the feasibility of the proposed constrained optimization problem, it suffices that the effective number of linear inequalities does not exceed $\sum_{m=0}^{M}\mathrm{rank}(\widehat{\Psi}(t_{m}))$, requiring that $\sum_{j=1}^{J}\mathrm{rank}(A_{j})$ remains within this bound. For large-sample analysis, we assume that $\mathrm{Var}(\mathbf{X}_{i})$ is positive definite and that $T_{i,\ell}$ has a continuous probability density function, as stated in condition (K1) in the Appendix. This implies that $\mathrm{rank}(\widehat{\Psi}(t_{m}))=2p$ with probability tending to one as $n\to\infty$ for every $m=0,1,\ldots,M$. Consequently, when the sample size $n$ is sufficiently large, we require that $\sum_{j=1}^{J}\mathrm{rank}(A_{j})\leq 2p(M+1)$. The examples provided in (2.5)–(2.13) satisfy this condition since $M_{j}\leq M$ for all $j\in J$ with $|J|\leq p$.
@@
Suppose the conditions in Proposition 2.2.1 hold. For any $M\geq 1$, let $\mathcal{G}_{M}=\{t_{m}\in[0,1]:m=0,1,\ldots,M\}$ be the finite grid used in (2.4) with $10=t_{0}<t_{1}<\cdots<t_{M}=$ 1. Suppose $\boldsymbol{\beta}_{0}$ is the true vector coefficient function satisfying the partial shape constraint of interest with the hypothesis $H_{0}$. We have
$\max_{t\in\mathcal{G}_{M}}\|\widehat{\boldsymbol{\beta}}(t)-\boldsymbol{\beta}_{0}(t)\|=O_{P}\bigg{(}h^{2}+\sqrt{\frac{\log n}{nh}}\bigg{)}.$ (2.14)
Remark 3.
Theorem 2.2.2 indicates that the shape-constrained kernel least square estimator achieves the same rate of convergence as the unconstrained estimator under the null hypothesis. Moreover, suppose the empirical distribution of $\mathcal{G}_{M}$ converges to the uniform distribution on $[0,1]$. Then, it can be verified that $\int_{0}^{1}\|\widehat{\boldsymbol{\beta}}(t)-\boldsymbol{\beta}_{0}(t)\|\,\mathrm{d}t=O_{P}(n^{-2/5}\sqrt{\log n})$, the same rate as the unconstrained estimator, similarly to Remark 1.